OpenStudy (anonymous):

Which of the following equations is written in slope-intercept form? 2x + 4y = 8 4y = -2x + 8 y = -one halfx + 2 x = -2y + 4

OpenStudy (anonymous):

Well, do you know what slope intercept form entails?

OpenStudy (anonymous):

slope-intercept form is y= mx +c i.e. the equation must be in form where y is on one side with a coefficient of one..

OpenStudy (anonymous):

It is usually written as y= mx+b, where m is the slope and b is a constant number

OpenStudy (anonymous):

Knowing this, we can start ruling out some answers

OpenStudy (anonymous):

2x+4y=8 is not the proper form, so we can rule it out. Same thing applies for x=-2y+4

OpenStudy (anonymous):

y = -one halfx + 2

OpenStudy (anonymous):

Yes, that is written in slope intercept form

OpenStudy (anonymous):

Do you understand why 4y= -2x+8 is not the correct answer?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

then there you go! Congrats!